# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v2

@@ II.4 Spectral scaling
where the energy transfer function $T(k)$ is defined as [35],
@@ VI End Matter
This section provides the foundation for the phase diagram in Table 1, through simulations that systematically vary the disorder distribution $\Delta\omega$, the source of the enstrophy injection, or frustration, $\omega_{i}$, in Eq. (2). First, however, we must outline a conjunction of two seemingly disparate turbulence theories.
@@ VI.1 Appendix A: Three thermodynamic phases
Figure 4: The three distributions in $\Delta\omega$ corresponding to the three thermodynamic phases in Table 1, showing only the positive component ($\omega_{i}$ is symmetric around 0). The upper limit on the $x$-axis corresponds to the maximum fluctuation $1/2L$.
Figure 5: Summary of a Phase 1 simulation, showing the evolution of static, phase-locked clumps. Upper row shows raw particle locations while bottom row shows topological charge with a colorscale.
@@
To conclude the theoretical conjunction, we show, in Figure 4, three distributions in $\omega_{i}$, corresponding to $\Delta\omega$ in Table 1. In our simulations, intrinsic frustration $\omega_{i}$ near the box-size, $1/2L$, acts simultaneously as $Z$ in Ref. [18] and $\Omega$ in Ref. [33]. Thus, if the frustration dispersion $\Delta\omega$ is too low and too narrow, the system generates active stress but lacks the transport coherence to move it, yielding an arrested glass vortex (Figure 3, Phase II in Table 1). Reproducing a ferromagnetic-like collapse; a cold, isotropic, uniform “frustration” ($\Delta\omega\rightarrow 0$) yields a collection of static clumps (Phase I in Table 1). Conversely, if the dispersion in $\omega_{i}$ is both sufficiently high and sufficiently broad (approaching the box-size), the system effectively facilitates the emergence of a space-filling Onsager dipole (Phase III in Table 1).
@@ VI.2 Appendix B: Derivation of advective transport
Thus, $\Delta\omega$ functions as the **agitation parameter** or granular temperature of the topological gas. This creates the central thermodynamic trade-off of the system:

### v3

@@ II.4 Spectral scaling
where the energy transfer function $T(k)$ is defined for a velocity field $\mathbf{v}$ as [35],
@@ VI End Matter
The End Matter provides the empirical and theoretical foundation for the phase diagram in Table 1, first through simulations that systematically vary the disorder distribution $\Delta\omega$, then through a rigorous theoretical derivation of the model’s hydrodynamic limit. First, however, we must outline a conjunction of two seemingly disparate turbulence theories.
@@ VI.1 Appendix A: Three thermodynamic phases
Figure 4: The three distributions in $\Delta\omega$ corresponding to the three thermodynamic phases in Table 1, showing only the positive component ($\omega_{i}$ is symmetric around 0). The lower limit on the $x$-axis is given by $v_{0}/2L$, the hypothetical “box-size fluctuation.”.
Figure 5: Summary of a Phase I simulation (top row) and a Phase II simulation emphasising the emergence of hard-core defects (bottom row). Topological charge is shown with a colorscale; timestamps are indicated.
@@
To conclude the theoretical conjunction, we show, in Figure 4, three distributions in $\omega_{i}$, corresponding to $\Delta\omega$ in Table 1. A matter of some importance, the intrinsic frustration $\omega_{i}$ near the box-size, $v_{0}/2L$, acts simultaneously as $Z$ in Ref. [18] and $\Omega$ in Ref. [33]. Thus, if the frustration dispersion $\Delta\omega$ is too low and too narrow, the system generates active stress but lacks the transport coherence to move it, yielding an arrested glass vortex (Figure 3, Phase II in Table 1). Reproducing a ferromagnetic-like collapse; a cold, isotropic, uniform “frustration” ($\Delta\omega\rightarrow 0$) yields a collection of static clumps (Phase I in Table 1). Conversely, if the dispersion in $\omega_{i}$ is both sufficiently high and sufficiently broad (near the box-size fluctuations, $v_{0}/2L$), the system effectively facilitates the emergence of a space-filling Onsager dipole (Phase III in Table 1).
@@ VI.2 Appendix B: Derivation of advective transport
Thus, $\Delta\omega$ functions as the agitation parameter or granular temperature of the topological gas. This creates the central thermodynamic trade-off of the system: